9. A horse canters away from its trainer in a straight line,moving 160m away in 17s....

9.
A horse canters away from its trainer in a straight line,moving 160m away in 17s. It then turns abruptly and gallops halfwayback in 6.8s. Calculate the (a) its average speed and (b) itsaverage velocity for the entire trip, using "away from the trainer"as the positive direction.
(A) Average Speed
Formula:
(B) Average Velocity
Answer #1
Concepts and reason

The concepts used to solve this problem are average speed and average velocity.

Initially, the expression for the average speed is used to calculate the average speed for the entire trip.

Finally, the expression for the average velocity is used to calculate the average velocity for the entire trip.

Fundamentals

Average speed is defined as the rate of change of distance with time.

The expression for the average speed is given by,

sˉ=dtotttot\bar s = \frac{{{d_{{\rm{tot}}}}}}{{{t_{{\rm{tot}}}}}}

Here, sˉ\bar s is the average speed, ttot{t_{{\rm{tot}}}} is the total time, and dtot{d_{{\rm{tot}}}} is the total distance.

Average velocity is defined as the rate of change of displacement with time.

The expression for the average velocity is given by,

vˉ=ΔxΔt\bar v = \frac{{\Delta x}}{{\Delta t}}

Here, Δx\Delta x represents the change in the displacement and Δt\Delta t represents the change in time.

(A)

The expression for the average speed sˉ\bar s in terms total distance dtot{d_{{\rm{tot}}}} and the time ttot{t_{{\rm{tot}}}} is,

sˉ=dtotttot\bar s = \frac{{{d_{{\rm{tot}}}}}}{{{t_{{\rm{tot}}}}}}

The total distance is given by,

dtot=d+d{d_{{\rm{tot}}}} = d + d'

Here, d is the distance covered in 17 s and dd' is the distance covered in 6.8 s.

Substitute 160 m for d and 80 m for dd' in above equation as follows:

dtot=160m+80m=240m\begin{array}{c}\\{d_{{\rm{tot}}}} = 160{\rm{ m}} + 80{\rm{ m}}\\\\ = 240{\rm{ m}}\\\end{array}

The total time is given by,

ttot=t+t{t_{{\rm{tot}}}} = t + t'

Here, t is the time taken cover 160 m and tt'is the time taken to cover half way distance.

Substitute 17 s for t and 6.8 s for tt' in above equation as follows:

ttot=17s+6.8s=23.8s\begin{array}{c}\\{t_{{\rm{tot}}}} = 17{\rm{ s}} + 6.8{\rm{ s}}\\\\ = 23.8{\rm{ s}}\\\end{array}

Substitute 240 m for dtot{d_{{\rm{tot}}}} and 23.8 s for ttot{t_{{\rm{tot}}}} in equation sˉ=dtotttot\bar s = \frac{{{d_{{\rm{tot}}}}}}{{{t_{{\rm{tot}}}}}} as follows:

sˉ=240m23.8s=10.08m/s\begin{array}{c}\\\bar s = \frac{{240{\rm{ m}}}}{{23.8{\rm{ s}}}}\\\\ = 10.08{\rm{ m/s}}\\\end{array}

(B)

The expression for the average velocity vˉ\bar vis,

vˉ=ΔxΔt\bar v = \frac{{\Delta x}}{{\Delta t}}

The change in displacement is given by,

Δx=x2x1\Delta x = {x_2} - {x_1}

Here, x1{x_1} and x2{x_2}are the initial and final positions.

Substitute 160 m for x2{x_2} and 80 m for x1{x_1}, in above equation as follows:

Δx=160m80m=80m\begin{array}{c}\\\Delta x = 160{\rm{ m}} - 80{\rm{ m}}\\\\ = 80{\rm{ m}}\\\end{array}

The change in time is given by,

Δt=t+t\Delta t = t + t'

Here, t is the time taken cover 160 m and tt'is the time taken to cover half way distance.

Substitute 17 s for t and 6.8 s for tt' in above equation as follows:

ttot=17s+6.8s=23.8s\begin{array}{c}\\{t_{{\rm{tot}}}} = 17{\rm{ s}} + 6.8{\rm{ s}}\\\\ = 23.8{\rm{ s}}\\\end{array}

Substitute 80 m for Δx\Delta x and 23.8 s for Δt\Delta t in equation vˉ=ΔxΔt\bar v = \frac{{\Delta x}}{{\Delta t}} as follows:

vˉ=80m23.8s=3.36m/s\begin{array}{c}\\\bar v = \frac{{80{\rm{ m}}}}{{23.8{\rm{ s}}}}\\\\ = 3.36{\rm{ m/s}}\\\end{array}

Ans: Part A

The average speed for the entire trip is 10.08m/s10.08{\rm{ m/s}}.